Kent Artists is a consignment gallery for local artists in Kent. “Consignment” means that the artists bring their original artwork to the gallery, the gallery displays, markets, and sells it, and the gallery shares the revenue from the sale with the artist. Right now, Kent Artists has a standard contract with its artists where 45% of the revenue from each sale goes to the artist and 55% goes to Kent Artists. Kent Artists uses their portion of the revenue to cover their fixed costs of maintaining, marketing, and staffing the gallery. At the moment, Kent Artists’ fixed costs total $20,000 per month.

1. What is Kent Artists’ contribution margin?

2. If Kent Artists sells a painting for $1500, what is their variable cost for that painting?

3. What is the unit contribution of the $1500 painting?

4. If Kent Artists makes $50,000 in revenue next month, what will their profit be for the month?

5. How much revenue does Kent Artists need to produce in a month to just cover their $20,000 fixed costs?
Answer #1

1. contribution margin = 55%

2. variable cost = 1500 * (1 - 0.55) = 675 (since variable cost ratio is 1- contribution margin ratio)

3. Contribution = 1500 * 55% = 825

4. Profit = Contribution - fixed cost

=(50,000 * 55%) - 20,000 = 7500

5. revenue to cover the fixed cost = fixed cost / contribution margin

= 20,000 / 55%

= 36,363.64